# Guiding Time-Varying Generative Models with Natural Gradients on Exponential Family Manifold

### v1

@@ Abstract
Optimising probabilistic models is a well-studied field in statistics. However, its connection with the training of generative models remains largely under-explored. In this paper, we show that the evolution of time-varying generative models can be projected onto an exponential family manifold, naturally creating a link between the parameters of a generative model and those of a probabilistic model. We then train the generative model by moving its projection on the manifold according to the natural gradient descent scheme. This approach also allows us to approximate the natural gradient of the KL divergence efficiently without relying on MCMC for intractable models. Furthermore, we propose particle versions of the algorithm, which feature closed-form update rules for any parametric model within the exponential family. Through toy and real-world experiments, we validate the effectiveness of the proposed algorithms.
@@ 1 Introduction
Modern generative models [Goodfellow et al., 2014, Ho et al., 2020, Song et al., 2021] have become indispensable tools in modern machine learning, achieving remarkable success in applications [Rombach et al., 2022, Gu et al., 2022, Li et al., 2019a, Tan et al., 2024]. These generative models are neural networks transforming a latent variable to a higher dimensional sample. They overcome classic restrictions imposed on probability density models, such as positivity, normalization, and encoding of conditional independence via factorization; thus, they can be designed freely to capture complex, intricate patterns from the high-dimensional data. Albeit these models produce highly realistic outputs, they can be hard to train. The model training requires massive data and maintains a delicate balance between “generators” and “discriminators” [Goodfellow et al., 2014, Arjovsky et al., 2017, Wang et al., 2023] or building effective “bridges” between the reference and target dataset [Song et al., 2021, Lipman et al., 2023, Liu et al., 2023, Bortoli et al., 2021], which are undesirable when samples are limited, or such bridges are difficult to design.
Parametric probabilistic models, particularly those in the exponential family [Casella and Berger, 2024, Wainwright et al., 2008], play a central role in modern day’s statistical inference, and have well-established theoretical framework and training algorithms. These models, characterised by their sufficient statistics and natural parameters, define a probability distributions manifold, the geometric structure of which inspired efficient optimisation methods such as natural gradient descent (NGD) Amari [1998], ensuring stable and efficient parameter estimation [Amari and Nagaoka, 2000]. However, while exponential family models are versatile in theory, their use may be limited in practice: the hand-crafted sufficient statistic may fail to capture complex relationships in data; more flexible sufficient statistics (such as neural nets) result in intractable likelihoods. Thus, parameter estimation that requires a likelihood, such as NGD, cannot be easily applied to fit the model.
@@
2 Backgrounds
In recent years, it has been recognized that there are many similarities between sampling and optimization [Wibisono, 2018, Arbel et al., 2019, Chewi, 2024, Cheng et al., 2018, He et al., 2025]. Encouraged by these results, we ask: Given a time-varying generative model $g_{t}$, where $t$ is time, can we move its output distribution $q_{g_{t}}$ toward an optimal distribution $q^{*}$ using NGD? In literature, generative models are often referred to as implicit models and generative distribution $q_{g_{t}}$ normally don’t have a parametric density, thus optimization designed for parametric densities cannot be directly applied.
Before stating our solutions, we first introduce a few ideas that are essential for developing the proposed algorithm.
@@ 2.1 Time-varying Exponential Family
(See, e.g., Section 3.4, Casella and Berger [2024]) A distribution is a member of the exponential family with sufficient statistic $T$ if and only if its density function $q({\boldsymbol{x}};{\boldsymbol{\theta}})$ can be expressed as:
@@
The parametric manifold of exponential family distributions with a sufficient statistic $T$ is:
@@
Equivalently, this can be expressed as the inner product between the rate of change of the natural parameter and the centered sufficient statistic (Proposition 3.1 in Williams et al. [2025]):
@@ 2.2 Natural Gradient Descent
Natural gradient descent is an effective optimization technique for parametric probabilistic models [Amari, 1998, Amari and Nagaoka, 2000] and has been widely applied in various fields [Bernacchia et al., 2018, Chen et al., 2024b].
@@
Although methods described in this paper are generic, we focus on KL divergence in this paper.
@@
Except for certain specific choices of $T$, $\nabla^{N}{\mathrm{KL}}$ does not admit a closed-form solution, as neither $\mathbb{E}_{q_{{\boldsymbol{\theta}}}}[T({\boldsymbol{x}})]$ nor $\mathrm{Cov}_{q_{{\boldsymbol{\theta}}}}[T({\boldsymbol{x}})]$ can be expressed in closed form for a general $T$. If we could sample from $q_{{\boldsymbol{\theta}}}$, these expectations and covariances could be approximated using Monte Carlo methods. However, generating samples from a complex distribution $q_{{\boldsymbol{\theta}}}$ itself remains a challenging problem.
@@ 2.3 Time-varying Generative Model
In recent years, there has been a growing trend of designing generative models as functions of time, in contrast to the classic generative models where the sample generating mechnism is independent of time. For example, the diffusion generative model [Song et al., 2021] can be interpreted as the solution to a Stochastic Differential Equation at time $t$, with initial samples drawn from a reference distribution. Similarly, the rectified flow generative model [Liu et al., 2023] can be viewed as the solution to an Ordinary Differential Equation at time $t$.
@@
A generative model is a data-generating mechanism defined as $X={\boldsymbol{g}}(Z;{\boldsymbol{w}})$, where $Z$ is a random variable sampled from a latent distribution $p_{Z}$. ${\boldsymbol{w}}\in\mathcal{W}$ are parameters of the generative model ${\boldsymbol{g}}$. A time-varying generative model is a generative model with time dependency, defined as $X_{t}={\boldsymbol{g}}(Z,t;{\boldsymbol{w}})$.
@@ 2.4 Time Score Matching
Given a parametric time score model ${v}({\boldsymbol{x}};t)$ and a time-varying sample $X_{t}\sim q_{t}$, the time score can be learned by the Time Score Matching (TSM) which minimizes the objective: $\int\mathbb{E}\Bigl{[}\lambda(t)\,\bigl{\|}s_{t}(X_{t})-{v}\bigl{(}X_{t};t\bigr{)}\bigr{\|}^{2}\Bigr{]}\,\mathrm{d}t,$ where $\lambda(t)$ is a weighting function.
@@ 3.1 Problem Formulation
Note that we do not want to train a generative model using GANs or diffusion models, as these models require a large number of samples and are hard to train. Instead, given a loss function $\mathcal{L}(p,q_{{\boldsymbol{\theta}}_{t}})$ that measures the difference between $p$ and a time-varying probabilistic model $q_{{\boldsymbol{\theta}}_{t}}$, we aim to guide the training of the generative model by minimizing the loss $\mathcal{L}$ of the probabilistic model $q_{{\boldsymbol{\theta}}_{t}}$ over time.
The key idea of this paper is to align the evolution of the generative model with $\nabla^{N}\mathcal{L}({\boldsymbol{\theta}})$ on the manifold $\mathcal{M}(T)$. Consequently, minimizing the loss for $q_{\boldsymbol{\theta}}$ simultaneously drives the generative model toward matching $p$.
@@ 3.2 Projecting the Change
At a fixed time $t_{0}$, we can “project” the time score $s_{t}$ onto the manifold $\mathcal{M}(T)$ by minimizing the squared difference between $s_{t}$ and the time score of an exponential family distribution, $\partial_{t}(\log q_{{\boldsymbol{\theta}}_{t}}),q_{{\boldsymbol{\theta}}_{t}}\in\mathcal{M}(T)$,
@@
which now depends on the parameter ${\boldsymbol{\delta}}$. One can view (4) as a local regression at the fixed time point $t_{0}$: $\lambda$ is a time smoothing kernel, Equation 4 finds the best score model that approximates the time-varying function $s_{t}({\boldsymbol{x}})$ at $t_{0}$.
The minimizer to the above objective is a vector in $\mathbb{R}^{d}$ that best describes the instantaneous change in the generative model. Moreover, we can show that
@@
The proof can be found in Appendix B. Since $X_{t}=g(Z,t;{\boldsymbol{w}})$, we can express ${\boldsymbol{\delta}}_{t_{0}}$ as a function of ${\boldsymbol{w}}$ using the reparameterization trick [Kingma and Welling, 2014].
@@
Different time scores can be mapped to the same projection ${\boldsymbol{\delta}}_{t_{0}}$, especially when the sufficient statistic $T$ is restrictive. However, if $T$ is chosen such that the exponential family is expressive enough, we expect that such information collapse could be avoided. The rigorous proof of this claim is left as a future work.
@@ 4 Special Case: Drift Model
We can prove that when using a model introduced in Definition 4.1, Section 3.2 has a limiting solution as $\sigma\to 0$, which eliminates the bias caused by the time-smoothing kernel $\lambda$.
@@
The proof can be found in Appendix C. Using this limiting solution, the objective of Equation 6 can be rewritten as:
@@ 4.1 Kernel NGD
Algorithm 1 Generative Model Training Guided by Natural Gradient Descent
@@
5:     Approximate ${\boldsymbol{\delta}}({\boldsymbol{w}})$ with $X_{t[i]}$ using Monte Carlo.
6:     ${\boldsymbol{w}}\leftarrow\mathop{\rm argmin}\limits_{{\boldsymbol{w}}}\|\nabla^{N}\mathcal{L}({\boldsymbol{\theta}}_{t[i]})-{\boldsymbol{\delta}}({\boldsymbol{w}})\|^{2}$
@@
The optimal drift function that minimizes Equation 8 can be found as
${\boldsymbol{h}}_{w^{*}}({\boldsymbol{x}})=\mathbb{E}\left[\nabla T(X_{t_{0}})\nabla k(X_{t_{0}},{\boldsymbol{x}})\right]^{\top}{\boldsymbol{\Gamma}}^{-1}\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})$ (9)
@@
The proof can be found in Appendix D. This result enables the direct calculation of particle updates without fitting a generative model first. This inspires us to build a particle evolution strategy guided by NGD: First, we sample $X_{0}$ from an initial distribution $q_{0}$, then we iteratively update each sample $X_{t_{0}}$ using the formula given by Theorem 4.3 until they converge. This algorithm is summarized in Algorithm 2 and we name it Kernel implicit NGD (KiNG).
@@
In Figure 3, we plot the trajectories of KiNG with different one-dimensional initial and target distributions. We choose $\mathcal{M}(T)$ to be a Gaussian manifold, i.e., $T(x)=[x,x^{2}]^{\top}$. Note that in the right plot, the particles do not converge to the bi-modal target distribution, since the movements of our particles are “guided” by the Gaussian manifold, and the best approximation is a Gaussian with a larger standard deviation. This example shows that the particles are indeed guided by the Gaussian manifold throughout the generative process. This phenomenon could be beneficial, if the target is to find the best approximation within a given family or we already have an informative probabilistic model (see Section 5.2).
This behaviour could be changed by replacing the Gaussian manifold with a more expressive manifold. In the left plot of Figure 4, we run similar experiments by letting $T$ be the Radial Basis Functions (RBFs) and we can see that indeed the particles bifurcate and converge to both modes.
@@ 4.2 Neural Tangent Kernel NGD
${\boldsymbol{h}}({\boldsymbol{x}}):=\nabla_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{x}};{\boldsymbol{\beta}}_{0}){\boldsymbol{w}},$
@@
The optimal drift that minimizes Equation 10 can be expressed using the neural tangent kernel [Jacot et al., 2018]:
@@
where ${\boldsymbol{K}}_{\mathrm{NTK}}$ is the matrix-valued neural tangent kernel, defined as ${\boldsymbol{K}}_{\mathrm{NTK}}({\boldsymbol{x}},{\boldsymbol{y}}):=\nabla_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{x}})\nabla^{\top}_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{y}})$. This result can be proven using the same technique described in Appendix D. In this paper, we use empirical and a finite-width NTK for simplicity, but NTKs that are infinitely wide can be efficiently computed using off-the-shelf package such as neural-tangents [Novak et al., 2020] for a variety of neural network architectures.
@@ 5 Experiments
In this section, we further validate our method on toy data and real-world data respectively. The summarized details of experiment setups can be found in Appendix G. In all experiments, $\mathcal{L}(p,q_{\boldsymbol{\theta}})=\mathrm{KL}[p,q_{\boldsymbol{\theta}}]$.
@@
In this experiment, we compare KiNG, ntKiNG, with reverse KL Wasserstein Gradient Flow (WGF) [Gao et al., 2019, Liu et al., 2024] and Maximum Mean Discrepancy flow (MMD flow) [Hagemann et al., 2024] on small datasets with different dimensions. See Appendix F for more explanations of these methods.
Since they all minimize different divergences, we measure their performance using MMD [Gretton et al., 2012] between a fresh batch of target samples and $X_{t}$.

### v2

@@ Abstract
Optimising probabilistic models is a well-studied field in statistics. However, its connection with the training of generative models remains largely under-explored. In this paper, we show that the evolution of time-varying generative models can be projected onto an exponential family manifold, naturally creating a link between the parameters of a generative model and those of a probabilistic model. We then train the generative model by moving its projection on the manifold according to the natural gradient descent scheme. This approach also allows us to efficiently approximate the natural gradient of the KL divergence without relying on MCMC for intractable models. Furthermore, we propose particle versions of the algorithm, which feature closed-form update rules for any parametric model within the exponential family. Through toy and real-world experiments, we validate the effectiveness of the proposed algorithms. The code of the proposed algorithms can be found at https://github.com/anewgithubname/iNGD.
@@ 1 Introduction
Modern generative models [Goodfellow et al., 2014, Ho et al., 2020, Song et al., 2021] have become indispensable tools in machine learning, achieving remarkable success in applications [Rombach et al., 2022, Gu et al., 2022, Li et al., 2019a, Tan et al., 2024]. These generative models are neural networks that transform a latent variable into a higher-dimensional sample. They overcome classic restrictions imposed on probability density models, such as positivity, normalisation, and encoding of conditional independence via factorisation; thus, they can be designed freely to capture complex, intricate patterns from the high-dimensional data. Although these models produce highly realistic outputs, they can be hard to train. Training them requires massive amounts of data and maintains a delicate balance between “generators” and “discriminators” [Goodfellow et al., 2014, Arjovsky et al., 2017, Wang et al., 2023] or building effective “bridges” between the reference dataset and target dataset [Song et al., 2021, Lipman et al., 2023, Liu et al., 2023, Bortoli et al., 2021] which can be difficult to design.
Parametric probabilistic models, particularly those in the exponential family [Casella and Berger, 2024, Wainwright et al., 2008], play a central role in modern statistical inference, and have well-established theoretical framework and training algorithms. These models, characterised by their sufficient statistics and natural parameters, define a probability distribution manifold, the geometric structure of which inspired efficient optimisation methods such as natural gradient descent (NGD) Amari [1998], ensuring stable and efficient parameter estimation [Amari and Nagaoka, 2000]. However, while exponential family models are versatile in theory, their use may be limited in practice: the hand-crafted sufficient statistic may fail to capture complex relationships in data; more flexible sufficient statistics (such as neural nets) result in intractable likelihoods. Thus, parameter estimation that requires a likelihood, such as NGD, cannot be easily applied to fit the model.
@@
2 Background
In recent years, it has been recognised that there are many similarities between sampling and optimisation [Wibisono, 2018, Arbel et al., 2019, Chewi, 2024, Cheng et al., 2018, He et al., 2025]. Encouraged by these results, we ask: Given a time-varying generative model $g_{t}$, where $t$ is time, can we move its output distribution $q_{g_{t}}$ toward an optimal distribution $q^{*}$ using NGD? In literature, generative models are sometimes referred to as implicit models and generative distribution $q_{g_{t}}$ may not have a parametric density [^1] Some generative models, such as normalizing flows [Rezende and Mohamed, 2015], neural ODEs [Chen et al., 2018], admit explicit parametric forms. , thus optimisation designed for parametric densities cannot be directly applied.
Before stating our solutions to this problem, we first introduce a few ideas that are essential for developing the proposed algorithm.
@@ 2.1 Time-varying Exponential Family
(See, e.g., Section 3.4, Casella and Berger [2024]) A distribution is a member of the exponential family with the sufficient statistic $T$ if and only if its density function $q({\boldsymbol{x}};{\boldsymbol{\theta}})$ can be expressed as:
@@
The parametric manifold of the exponential family distributions with a sufficient statistic $T$ is:
@@
Equivalently, this can be expressed as the inner product of the natural parameter’s rate of change and the centred sufficient statistic (Proposition 3.1 in Williams et al. [2025]):
@@ 2.2 Natural Gradient Descent
The natural gradient descent is an effective optimisation technique for parametric probabilistic models [Amari, 1998, Amari and Nagaoka, 2000] and has been widely applied in various fields [Bernacchia et al., 2018, Chen et al., 2024b].
@@
Although methods described in this paper are generic for different $\mathcal{L}$, we focus on the KL divergence in this paper.
@@
Except for certain specific choices of $T$, $\nabla^{N}{\mathrm{KL}}$ does not admit a closed-form expression, as neither $\mathbb{E}_{q_{{\boldsymbol{\theta}}}}[T({\boldsymbol{x}})]$ nor $\mathrm{Cov}_{q_{{\boldsymbol{\theta}}}}[T({\boldsymbol{x}})]$ can be expressed in closed form for a general $T$. If we could sample from $q_{{\boldsymbol{\theta}}}$, these expectations and covariances could be approximated using Monte Carlo methods. However, generating samples from a complex distribution $q_{{\boldsymbol{\theta}}}$ itself remains a challenging problem.
@@ 2.3 Time-varying Generative Model
In recent years, there has been a growing trend of designing generative models as functions of time, in contrast to the classic generative models where the sample generating mechanism is independent of time. For example, the diffusion generative model [Song et al., 2021] can be interpreted as the solution to a Stochastic Differential Equation at time $t$, with initial samples drawn from a reference distribution. Similarly, the rectified flow generative model [Liu et al., 2023] can be viewed as the solution to an Ordinary Differential Equation at time $t$.
@@
A generative model is a data-generating mechanism defined as $X={\boldsymbol{g}}(Z;{\boldsymbol{w}})$, where $Z$ is a random variable sampled from a latent distribution $p_{Z}$. ${\boldsymbol{w}}\in\mathcal{W}$ are parameters of the generative model ${\boldsymbol{g}}$. A time-varying generative model is defined as $X_{t}={\boldsymbol{g}}(Z,t;{\boldsymbol{w}})$, i.e., a generative model whose output depends explicitly on $t$.
@@ 2.4 Time Score Matching
Given a parametric time score model ${v}({\boldsymbol{x}};t)$ and a time-varying sample $X_{t}\sim q_{t}$, the time score can be learned by the Time Score Matching (TSM) which minimises the objective: $\int\mathbb{E}\Bigl{[}\lambda(t)\,\bigl{\|}s_{t}(X_{t})-{v}\bigl{(}X_{t};t\bigr{)}\bigr{\|}^{2}\Bigr{]}\,\mathrm{d}t,$ where $\lambda(t)$ is a weighting function.
@@ 3.1 Problem Formulation
We want to avoid training generative models using GANs or diffusion models, as these models require a large number of samples and are hard to train. Instead, given a loss function $\mathcal{L}(p,q_{{\boldsymbol{\theta}}_{t}})$ that measures the difference between $p$ and a time-varying probabilistic model $q_{{\boldsymbol{\theta}}_{t}}$, we aim to guide the training of the generative model by minimising the loss $\mathcal{L}$ of the probabilistic model $q_{{\boldsymbol{\theta}}_{t}}$ over time.
The key idea of this paper is to align the evolution of the generative model with $\nabla^{N}\mathcal{L}({\boldsymbol{\theta}})$ on the manifold $\mathcal{M}(T)$. Consequently, minimising the loss for $q_{\boldsymbol{\theta}}$ simultaneously drives the generative model toward matching $p$.
@@ 3.2 Projecting the Change
At a fixed time $t_{0}$, we can “project” the time score $s_{t}$ onto the manifold $\mathcal{M}(T)$ by minimising the squared difference between $s_{t}$ and the time score of an exponential family distribution, $\partial_{t}(\log q_{{\boldsymbol{\theta}}_{t}}),q_{{\boldsymbol{\theta}}_{t}}\in\mathcal{M}(T)$,
@@
which depends on the parameter ${\boldsymbol{\delta}}$. We use linear-in-time model to simplify the optimisation problem. More model choices of ${\boldsymbol{\theta}}(t)$ could be found in [Williams et al., 2025]. (4) can be viewed as a local regression at the fixed time point $t_{0}$: $\lambda$ is a time smoothing kernel, Equation 4 finds the best score model that approximates $s_{t}({\boldsymbol{x}})$ at $t_{0}$.
The minimiser to the above objective is a vector in $\mathbb{R}^{d}$ that best describes the instantaneous change in the generative model. Moreover, we can show that
@@
The proof can be found in Appendix A. Since $X_{t}=g(Z,t;{\boldsymbol{w}})$, we can express ${\boldsymbol{\delta}}_{t_{0}}$ as a function of ${\boldsymbol{w}}$ using the reparameterization trick [Kingma and Welling, 2014].
@@
[…]
@@ 4 Special Case: Drift Model
We can prove that when using a model introduced in Definition 4.1, Section 3.2 has a limiting solution as $\sigma\to 0$, which eliminates the bias caused by the smoothing kernel $\lambda$.
@@
The proof can be found in Appendix B. Using this limiting solution, the objective of Equation 6 can be rewritten as:
@@ 4.1 Kernel NGD
Algorithm 1 Generative Model Training Guided by the Natural Gradient Descent
@@
5:     Approximate ${\boldsymbol{\delta}}_{t[i]}({\boldsymbol{w}})$ with $X_{t[i]}$ using Monte Carlo.
6:     ${\boldsymbol{w}}\leftarrow\mathop{\rm argmin}\limits_{{\boldsymbol{w}}}\|\nabla^{N}\mathcal{L}({\boldsymbol{\theta}}_{t[i]})-{\boldsymbol{\delta}}_{t[i]}({\boldsymbol{w}})\|^{2}$
@@
The optimal drift function that minimises Equation 8 can be found as
${\boldsymbol{h}}_{w^{*}}({\boldsymbol{x}})=\mathbb{E}\left[\nabla T(X_{t_{0}})\nabla\nabla k(X_{t_{0}},{\boldsymbol{x}})\right]^{\top}{\boldsymbol{\Gamma}}^{-1}\nabla\mathcal{L}({\boldsymbol{\theta}}_{t_{0}})$ (9)
@@
The proof can be found in Appendix C. This result enables the direct calculation of particle updates without fitting a generative model first. This inspires us to build a particle evolution strategy guided by NGD: First, we sample $X_{0}$ from an initial distribution $q_{0}$, then we iteratively update each sample $X_{t_{0}}$ using the formula given by Theorem 4.3 until they converge. This algorithm is summarized in Algorithm 2 and we name it Kernel implicit NGD (KiNG).
@@
In Figure 3, we plot the trajectories of KiNG with different one-dimensional initial and target distributions. We choose $\mathcal{M}(T)$ to be a Gaussian manifold, i.e., $T(x)=[x,x^{2}]^{\top}$. In the right plot, the particles do not converge to the bi-modal target distribution, since the movements of our particles are “guided” by the Gaussian manifold, and the best approximation is a Gaussian with a larger standard deviation. This example shows that the particles are indeed guided by the Gaussian manifold throughout the generative process. This phenomenon could be beneficial, if the target is to find the best approximation within a given family or we already have an informative probabilistic model (see Section 5.2).
This behaviour can be changed by replacing the Gaussian manifold with a more expressive manifold. In the left plot of Figure 4, we run similar experiments by letting $T$ be the Radial Basis Functions (RBFs) and we can see that indeed the particles bifurcate and converge to both modes.
@@ 4.2 Neural Tangent Kernel NGD
${\boldsymbol{h}}({\boldsymbol{x}};{\boldsymbol{w}}):=\nabla_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{x}};{\boldsymbol{\beta}}_{0}){\boldsymbol{w}},$
@@
The optimal drift that minimises Equation 10 can be expressed using the neural tangent kernel [Jacot et al., 2018]:
@@
where ${\boldsymbol{K}}_{\mathrm{NTK}}$ is the matrix-valued neural tangent kernel, defined as ${\boldsymbol{K}}_{\mathrm{NTK}}({\boldsymbol{x}},{\boldsymbol{y}}):=\nabla_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{x}})\nabla^{\top}_{\boldsymbol{\beta}}{\boldsymbol{\phi}}({\boldsymbol{y}})$. This result can be proven using the same technique described in Appendix C. In this paper, we use empirical and a finite-width NTK for simplicity, but NTKs that are infinitely wide can be efficiently computed using off-the-shelf package such as neural-tangents [Novak et al., 2020] for a variety of neural network architectures.
@@ 5 Experiments
In this section, we further validate our method on toy data and real-world data respectively. The summarized details of experiment setups can be found in Appendix F. In all experiments, $\mathcal{L}(p,q_{\boldsymbol{\theta}})=\mathrm{KL}[p,q_{\boldsymbol{\theta}}]$.
@@
In this experiment, we compare KiNG, ntKiNG, with reverse KL Wasserstein Gradient Flow (WGF) [Gao et al., 2019, Liu et al., 2024] and Maximum Mean Discrepancy flow (MMD flow) [Hagemann et al., 2024] on small datasets with different dimensions. See Appendix E for more explanations of these methods.
Since they all minimise different divergences, we measure their performance using MMD [Gretton et al., 2012] between a fresh batch of target samples and $X_{t}$.